Advanced Company reports the following information for the current year. All beginning inventory amounts equaled $0 this year. Units produced this year 33,000 units Units sold this year 19,800 units Direct materials $ 17 per unit Direct labor $ 19 per unit Variable overhead $ 3 per unit Fixed overhead $ 165,000 in total

Given Advanced Company's data, compute cost per unit of finished goods under variable costing.

Given Advanced Company's data, compute cost per unit of finished goods under absorption costing.

Given Advanced Company's data, compute cost of finished goods in inventory under variable costing.

Given Advanced Company's data, and the knowledge that the product is sold for $64 per unit and operating expenses are $280,000, compute the net income under absorption costing.

Given Advanced Company's data, and the knowledge that the product is sold for $64 per unit and operating expenses are $280,000, compute the net income under variable costing.
Answer #1
Advance Company
Absorption Costing Method
Cost per unit of Finished goods
Direct Material Per unit 17
Direct Labour per unit 19
Varaible overheads per unit 3
Fixed overheads per unit (165000/33000) 5
Cost per unit of Finished goods 44
Cost per unit of Finished goods under Absorption costing
Cost per unit of Finished goods 44
Number of units in ending finished goods 13200
Cost of goods sold
Cost per unit of Finished goods 44
Number of units sold 19800
Net Income under Absorption costing
Sales 1267200
Less: Cost of goods sold 871200
Gross profit 396000
Less: Operating Expense 280000
Net Income under Absorption costing 116000
Varaible Costing
Cost per unit of Finished goods under Variable costing
Cost per unit of Finished goods= Direct materials+ Direct Labour +Varaiable overheads
Cost per unit of Finished goods= 17+19+3= 39
Net Income under Variable costing
Particulars Amount $ Amount $
A)Sales   (19800 units*$64) 1267200
Less: Varaible Cost
Operating Expenses 280000
Direct Materials 561000
Direct Labour 627000
Variable Overheads 99000
Total Varaible Cost 1567000
Less:Ending Finished goods Inventory -514800
B) 1052200
Contribution Margin (A-B) 215000
Less: Fixed Overheads 165000
Net Income 50000
Direct Materials 33000*17= 561000
Direct Labour 33000*19= 672000
Variable Overheads 33000*3= 99000
Ending Finished goods inventory
Closing stock* total variable cost
(33000-19800)*(17+19+3)= 514800